OpenStudy (anonymous):

NEVER combine terms across an equal sign. You may do anything you want to an equation, as long as you do it to _______ sides of the equation

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

If I have 2=2 I'll add 3 both the sides 2+3=2+3 5=5 So that the equation still remains the same, isn't it?

OpenStudy (anonymous):

Yes

OpenStudy (ash2326):

If I add 3 to only one side \[5\ne 2\] So I have to do the addition to ?

OpenStudy (anonymous):

Yes to get 5 right ?

OpenStudy (ash2326):

nope to make the equation remain unchanged. I have to perform the operation ..... the sides

OpenStudy (anonymous):

I just got conused ash :c

OpenStudy (ash2326):

oh when I added 3 to both the sides, our equation held true. But when I added 3 to one side, our equation was changed and became false. Did you follow this part

OpenStudy (anonymous):

Oh yes okay .. So what would I fill in my blank? Both sides?

OpenStudy (ash2326):

yes, right, when we do something both the sides. Our equation doesn;t change and help us get the solution

OpenStudy (anonymous):

Yay thank you Ash (: